please help me to find the value of "a" in the equation (5)/(a+3)=(3)/(a-2)
Answer by fractalier(6550)

Let us cross-multiply and solve...we have
(5)/(a+3)=(3)/(a-2)
and then
5(a-2) = 3(a+3) expand
5a - 10 = 3a + 9 subtract 3a
2a - 10 = 9 add 10
2a = 19 divide by 2
a = 19/2